OpenStudy (anonymous):

Solve the equation. √3x - 5 = 4 Solve the equation. 4√x - 5 - 8 = 12 Solve the equation. Check both solutions and only write the real solution. √6m -5 = m

Parth (parthkohli):

1) sqrt3x = 9 3x = 81 x = 27 2) 4sqrtx = 25 sqrt x = 25/4 x = 625/16 Can't figure 3

OpenStudy (earthcitizen):

Recall,\[\sqrt[n]{x} = a \therefore x = a ^{n}\]